### v2

# Neuro-Informed Joint Learning Enhances Cognitive Workload Decoding in Portable BCIs

Abstract
Portable and wearable consumer-grade electroencephalography (EEG) devices, like Muse headbands, offer unprecedented mobility for daily brain-computer interface (BCI) applications, including cognitive load detection. However, the exacerbated non-stationarity in portable EEG signals constrains data fidelity and decoding accuracy, creating a fundamental trade-off between portability and performance. To mitigate such limitation, we propose MuseCogNet (Muse-based Cognitive Network), a unified joint learning framework integrating self-supervised and supervised training paradigms. In particular, we introduce an EEG-grounded self-supervised reconstruction loss based on average pooling to capture robust neurophysiological patterns, while cross-entropy loss refines task-specific cognitive discriminants. This joint learning framework resembles the bottom-up and top-down attention in humans, enabling MuseCogNet to significantly outperform state-of-the-art methods on a publicly available Muse dataset and establish an implementable pathway for neurocognitive monitoring in ecological settings.
Index Terms:
BCI, EEG, Self-Supervised Learning, Cognitive Load, Wearable, Human Attention
I Introduction
Overloaded cognitive states in critical scenarios like driving are major contributors to car accidents [1], necessitating real-time cognitive workload detection. EEG-based BCIs offer promise due to non-invasiveness and temporal resolution. Yet researches based on medical-grade EEG devices [2, 3] face deployment barriers: their complex setups require expertise and continuous monitoring. Consumer-grade devices like Muse headbands provide accessibility but introduce amplified non-stationarity and signal variability from mobile operation. This demands robust algorithms to identify neural signatures amid environmental noise to enhance decoding accuracy. Self-supervised learning (SSL) provides a principled framework for learning noise-invariant representations directly from unlabeled EEG. By designing pretext tasks that exploit inherent electrophysiological pattern [4], SSL models can discover neural signatures resilient to environmental interference. However, traditional voltage reconstruction objective has been reported to underperform due to inherent signal instability [5]. We propose that reconstructing temporally aggregated (average-pooled) representations is more physiologically meaningful. Specifically, by forcing the model to reconstruct the average-pooled temporal segments, we argue that the model can integrate neuroinformative insights while attenuating transient noise, thus capturing sustained neural processes underlying cognitive states. For downstream classification, we integrate supervised learning (SL) to complement SSL’s physiological grounding. While SSL reveals fundamental neural dynamics, SL optimizes discriminative boundaries by amplifying task-specific domain-relevant representations. The SSL-SL synergy forms a computational duality: SSL’s bottom-up feature extraction and SL’s top-down decision refinement jointly emulate the human neurocognitive processing stream, where sensory-level representations are progressively modulated by task-specific demands.
The main contribution of this paper is summarized below:
•We propose a novel SSL reconstruction objective based on average-pooled EEG, adapting to EEG’s intrinsic non-stationary nature.
•We implement the first joint SSL-SL framework deployed on wearable mobile EEG headsets (Muse), establishing a portable paradigm for ecological cognitive load monitoring.
II Methodology
Encoder: We deploy parallel convolutions with increasing kernel lengths $(1\times 16)$, $(1\times 32)$, $(1\times 64)$, $(1\times 128)$ to capture spectral features $>$16Hz, 8Hz, 4Hz, and 2Hz (since sampling rate is 256Hz). Concatenated features undergo batch normalization and ELU activation, followed by dual temporal pooling (variance + average) [6]. Pooled streams merge via $(2\times 1)$ convolution to form latent representations.
Dual-Path Decoding: - SSL Path: Reconstructs average-pooled signals via two linear layers - SL Path: Projects latent features to logits via 1D convolution
Losses: Losses are tailored individually for SSL and SL components. The self-supervised reconstruction loss $\mathcal{L}_{\text{pool}}$ measures the difference between true average pooled EEG $\bm{X}_{\text{avg\_pooled}}$ and reconstructed average pooled EEG $\hat{\bm{X}}_{\text{avg\_pooled}}$ to force model to learn temporally robust features, easing the adversial effect of non-stationarity in EEG. For a given EEG segment, the average-pooled loss is defined as below:
$\mathcal{L}_{\text{pool}}=\frac{1}{C\times T^{\prime}}\sum_{c=1}^{C}\sum_{t=1}^{T^{\prime}}\left(X_{\text{avg\_pooled}}^{(c,t)}-\hat{X}_{\text{avg\_pooled}}^{(c,t)}\right)^{2}$ (1)
where $C$ is the number of channels, $T^{\prime}$ is the pooled dimension of EEG signals.
The supervised cross-entropy loss refines the features into task-specific representations. For any given EEG segment, this is defined by:
$\mathcal{L}_{\text{CE}}=-\sum_{k=1}^{K}y_{k}\log\left(p_{k}\right)$ (2)
where $K$ indicates the number of possible categories (three in this case, namely low, medium or high level of cognitive load), and $p$ is computed via softmax normalization of the SL decoded logits.
The overall loss for a given sample is calculated as follows:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{pool}}$ (3)
where $\lambda$ is the SSL loss strength, which is set to 0.5 empirically.
III Experiment and Analysis
III-A Dataset and Experimental Setup
We utilize the CL-Drive dataset [7] captured using Muse S headbands during driving simulations. Reported cognitive load scores from 21 participants are categorized into 3 levels (low/medium/high) for each 10-second intervals. To maximize label-temporal relevance, models received the final 2-second segment of each interval.
EEG data is bandpass-filtered (0.1-75 Hz) with 60Hz notch filtering. Training employed AdamW (lr=0.001) with 100 epochs and batch size 64. Evaluation follows leave-one-subject-out (LOSO) protocol: each participant served as test set once, mimicking real-world new-user scenarios. Reported metrics are averaged performance across all test subjects.
TABLE I: Comparison of LOSO Performance across Models
Models EEGNet Deep ConvNet ResNet MuseCogNet
Accuracy (%) $57.07$ $57.07$ $58.41$ 62.68
F1 score (%) $52.09$ $51.25$ $51.61$ 57.24
TABLE II: Ablation Results
Models W/O SSL MuseCogNet
Accuracy (%) $60.77$ 62.68
F1 score (%) $54.60$ 57.24
III-B Analysis and Discussion
The comparative analysis reveals distinct performance advantages of our proposed MuseCogNet. As seen in Table I, established models like EEGNet [8] and Deep ConvNet [9] demonstrate competent baseline performance (Accuracy: 57.07%, F1: 51.25-52.09%), validating their status as canonical architectures for EEG decoding. Meanwhile, our implementation of ResNet achieves 58.41% accuracy, closely matching the original CL-Drive benchmark(58.13%) [7], confirming proper experimental replication. Critically, MuseCogNet achieves superior performance of 62.68% accuracy and 57.24% F1, representing absolute gains of 4.27% accuracy and 5.63% F1 over the strongest baseline (ResNet). We attribute this significant margin to our novel joint learning paradigm enhanced by average-pooled reconstruction objective, where the SSL component learns steady neurodynamic patterns that conventional supervised CNNs fail to capture. The ablation study (Table II) provides definitive evidence: Removing SSL components causes performance degradation of 1.91% accuracy and 2.64% F1, confirming that our average-pooled reconstruction objective contributes substantially to feature robustness.
Overall, MuseCogNet enables reliable cognitive load monitoring on portable EEG by leveraging SSL-SL synergy—where SSL’s bottom-up feature extraction and SL’s top-down discriminant refinement exhibit functional similarities to cortical sensory-cognitive processing hierarchies. This joint processing architecture not only achieves superior performance but also suggests intriguing parallels to human attentional mechanisms worthy of future exploration.